OpenStudy (anonymous):

A boy drops a 2.50 kg rock down a 250 m well and listens for the echo. The air temperature is 0°C. How long after the stone is dropped will the boy hear the echo? The speed of sound through air at 0°C is v = 331 m/s. ***Not sure how to solve this? Thank you!! :)

OpenStudy (michele_laino):

the requested time is the time needed to the rock for reaching the bottom of the well plus the time needed to sound to reach the boy.

OpenStudy (anonymous):

oh!

OpenStudy (michele_laino):

now the time needed to the rock to reaching the bottom of the well is given by the subsequent formula: \[\Large t = \sqrt {\frac{{2h}}{g}} \] where g is the gravity, namely g=9.81 m/sec^2, and h=250 meters. So what is: \[\Large t = \sqrt {\frac{{2h}}{g}} = \sqrt {\frac{{2 \times 250}}{{9.81}}} = ...?\]

OpenStudy (anonymous):

okay, so then we get 7.142857143? is that the final answer? :/ it takes about 7 seconds?

OpenStudy (michele_laino):

the time needed to sound for reaching the boy is: t=250/331=0.755 seconds, so the requested time, is: 7.14+0.755=7.895 seconds

OpenStudy (michele_laino):